OpenStudy (shelby1290):

Write an equation in STANDARD FORM for the line that: a) passes through the point (4,5) and has slope m=2/3 b) passes through the points (1,-6) and (-3,-9)

OpenStudy (anonymous):

before we start, you have to learn some terms: standard form for a line is \(Ax + By = C\) vertex form of a line is \(y=mx+b\) , where m is the slope and b is the y-int equation of a line is \((y-y_1)=m(x-x_1)\), where \((x_1, y_1)\) is any point in the line and m is the slope.. let's do letter a , first so you are given the slope and a point use the equation of the line \((y-5)=(\frac{2}{3})(x-4)\) can you simplify this equation?

OpenStudy (shelby1290):

y=2/3(x-4)+5 sorry im not really good at this unit

OpenStudy (shelby1290):

@Chibi_Robo3 ^^

OpenStudy (anonymous):

yeah, that's great :) but since we are looking for STANDARD FORM we just have to leave it like this: \(y−5=\frac{2}{3}x−\frac{2}{3}(4)\\y-5=\frac{2}{3}x−\frac{8}{3}\\y=\frac{2}{3}x−\frac{8}{3}+5\\y=\frac{2}{3}x+\frac{7}{3}\) arranging and simplifying this equation, you'll get: \(3y=2x+7\) -> multiply both sides by 3 to remove fractions \(2x-3y=-7\) -> standard form \(ax + by= c \) ☺

OpenStudy (shelby1290):

oh okay so 2x-3y=-7 is the answer for part A ?

OpenStudy (anonymous):

for the next one you have to find the slope (m) using the slope formula: \( \LARGE m= \frac{Δy}{Δx}=\frac{y_A-y_B}{x_A-x_B} \) where \((x_A, y_A)=(1,-6)\) and \((x_B, y_B)= (-3,-9)\) can you solve for the slope? after you find the slope, do the same process as I did before for part A and yes:)

OpenStudy (shelby1290):

okay so i'd be doing the y2-y1/ x2-x1 strategy right?

OpenStudy (shelby1290):

so -9-(-6) / -3-1

OpenStudy (shelby1290):

= -9-(-6)=-3

OpenStudy (shelby1290):

and -3-1= -4

OpenStudy (shelby1290):

so m= -3/-4

OpenStudy (shelby1290):

yea that's all i know for that part because i'm not sure if i have to find B or not

OpenStudy (anonymous):

you don't need to find b , since we are looking for the standard form, you can just use the equation of a line

OpenStudy (shelby1290):

so would i do this? : y=mx+b y=-3/-4x+b (1,-6) -6=-3/-4x+1+b b= -6-(-3/-4)(1)

OpenStudy (shelby1290):

i get stuck on that part because i dont know how to subtract and multiply for that part

OpenStudy (shelby1290):

and im not sure if im following your steps properly lol

OpenStudy (anonymous):

for m , you got it right m=3/4 hmm..it's the same thing but i guess you understand the other one, so let's stick with your method:) what did you get for b then? (1,-6) \(y=\frac{3}{4}x+b\\-6=\frac{3}{4}(1)+b\\b=\frac{3}{4}+6\\b=?\)

OpenStudy (shelby1290):

b= 9/4

OpenStudy (shelby1290):

right?

OpenStudy (anonymous):

i got a different answer for b

OpenStudy (shelby1290):

oh then i probably got it wrong...what was your answer

OpenStudy (anonymous):

27/4

OpenStudy (shelby1290):

oh okay then i'll just go with your answer as b

OpenStudy (shelby1290):

but how did you get 27/4

OpenStudy (shelby1290):

like was it using your way?

OpenStudy (shelby1290):

@Chibi_Robo3 ^

OpenStudy (shelby1290):

oh okay i believe i understand what you did.... so you multiplied 6x4 and then added 3 and got 27/4

OpenStudy (shelby1290):

now that i understand...how can i put that into standard form?

OpenStudy (anonymous):

you have \(y=\frac{3}{4}x+\frac{27}{4}\) to change this vertex form to standard form multiply both sides by 4, to remove fractions, what will you get?

OpenStudy (shelby1290):

how do i do that ?

OpenStudy (anonymous):

just multiply each term by 4